If the above problems is carried out as indicated, but with the additional provision that the process is also performed at a constant temperature of 327.9 F, calculate the change of entropy for the steam during the process, Btu/b.°R.
1
Expert's answer
2020-03-19T11:29:22-0400

So, "\Delta Q=889.2" "Btu\/lb"


The change of entropy ("T=327.9 F=787.5 R=const" )


"\Delta S=\int{\frac{\delta Q}{T}}=\frac{889.2}{787.5}=1.129" "\frac{Btu}{lb\cdot °R}" Answer